OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2

OpenStudy (anonymous):

in the end, I solved for y and then y = 2x-2, then I plugged it in the other one and x turns out to be equal to 0.18181818818818188 AND I'M LIKE WHATTTT

OpenStudy (anonymous):

combine the like turms

OpenStudy (anonymous):

okay I already did that... read what I did before

OpenStudy (anonymous):

LOL, No. You don't solve them. Plot them on a graph.

OpenStudy (anonymous):

what u ahev to plot them?

OpenStudy (anonymous):

yuppers both are correct once you combine like terms that is as far as you go

OpenStudy (anonymous):

1, Discriminant = 49 - (4*5*3) = -11 2, nemaproblema: x² - 7/5x = -3/5 (x - 7/10)² - 0.49 = -0.6 (x - 7/10)² = -0.11 Both indicates there is no solution.

OpenStudy (anonymous):

cause the main question is How many solutions are there to the following system of equations? a. 0 b. 1 c. 2 d. infinately many

OpenStudy (anonymous):

can I give two pple medals?

OpenStudy (anonymous):

A there r no solutions!

OpenStudy (anonymous):

no just one :(

OpenStudy (anonymous):

Put the system into slope intercept form. and plot it on the graph. You can see how many solutions there are from how many times they intercept.

OpenStudy (anonymous):

I gave u the answer there are no solutions!!!!!!!!!!

OpenStudy (anonymous):

brianna I know where you got that answer but it's not. I looked it up too but they don't have the same question.

OpenStudy (anonymous):

this is true you got an infinite number which usually means that there is no true answer!

OpenStudy (anonymous):

alright imma graph it

OpenStudy (anonymous):

yes that will give you an ACCURATE answer

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100115131935AAGxXYR this is not it

OpenStudy (anonymous):

ok its indefanitly many!

OpenStudy (anonymous):

There is only 1 solution. Meaning the two lines only intercept once.

OpenStudy (anonymous):

but when u graph they can only intercept one or non times!

OpenStudy (anonymous):

yes they intercept once.

OpenStudy (anonymous):

False, they can intercept infinite, zero, or one time.

OpenStudy (anonymous):

Infinite = Same Line Zero = Parallel Lines One = Criss Cross Lines.

OpenStudy (anonymous):

Stee, I am giving you a picture of the graph to help you. :)

OpenStudy (anonymous):

Here you go!

OpenStudy (anonymous):

ahhh I see thanks!!!